OpenStudy (anonymous):

What is the diameter of a circle whose circumference is 89.49 centimeters? Use 3.14 for π. 14.25 cm 5.34 cm 29.83 cm 28.5 cm

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Circumference = 2*π*r diameter = r*2 π = 3.14 Circumference = 89.46 diameter*3.14 = 89.46 diameter = 89.46/3.14 diameter = 28.49..... round to 28.5

OpenStudy (anonymous):

Source : google search :P

OpenStudy (anonymous):

lol

OpenStudy (cutebunny12):

this helped me thx